1.what volume of solution will contain 15g of NaCl if the solution concentration is 0.75

A. 0.33L B, 0.34L
C.0.35L. D,0.36L

2.how many kilograms of a 6.8% KC2 H3 O3 solution will contain 5.3 of KC2 H3 O2

A,7.5kg. B,7.6kg
C,7.7kg. D,7.8kg

3. 12.6g of NaCl are dissolved in water making 344ml of solution calculate the molar concentration

A,0.627M B,0.628M
C,0.629M D,0.630M
1
Expert's answer
2018-06-22T05:01:36-0400
1. n(NaCl) = m/M = 15g/58.44g/mol = 0.26 mol
V(solution) = n/C = 0.26mol/0.75mol/L = 0.35 L
C. 0.35 L
2. D. 7.8 kg
3. n(NaCl) = 12.6g/58.44g/mol = 0.216 mol
C = n/V = 0.216mol/0.344L = 0.628 mol/L
B. 0.628 M